# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ 3.1   Benefit matrix and the stability of symmetry
Let $\omega^{*}$ be the maximum possible utility gain from migration for a relocating agent upon an infinitesimal population deviation at $\bar{\bm{x}}$. Then, $\omega^{*}$ equals the largest eigenvalue of the benefit matrix $\mathbf{V}=\frac{\bar{x}}{\bar{v}}[\frac{\partial{v_{i}}}{\partial{x_{j}}}(\bar{\bm{x}})]_{i,j\in\mathcal{I}}$, and $\omega^{*}$ is achieved by deviation toward the associated eigenvector direction $\bm{z}^{*}$. ∎
Proof.
See Section A.1. ∎
The stability of $\bar{\bm{x}}$ depends on the sign of $\omega^{*}$. If $\omega^{*}<0$, any deviation lowers the utility of movers, implying local stability. Conversely, if $\omega^{*}>0$, movers benefit from deviating, triggering further migration. As a result, $\bar{\bm{x}}$ becomes unstable, and population starts to agglomerate along the direction $\bm{z}^{*}$.
Thus, we need to understand the eigenvalues of the benefit matrix $\mathbf{V}$. Our setting allows a key simplification on this regard:
Observation 2.
Under Assumption C, for each model (i.e., for each specific indirect utility function $\bm{v}$), the structure of the benefit matrix $\mathbf{V}$ remains unchanged from its two-region version. Specifically, if $\mathbf{V}=\Omega(\mathbf{D})$ with some matrix function $\Omega$ in the symmetric two-region setting, then we can just replace $\mathbf{D}$ with its many-region version to obtain the many-region version of $\mathbf{V}$. ∎
This allows a transparent approach for the stability analysis of $\bar{\bm{x}}$. For example, the formula $\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D}$ for the Beckmann model remains valid in the $N$-region economy, provided that $\mathbf{I}$ and $\mathbf{D}$ are replaced with their $N$-region counterparts. Then, the $k$th eigenvalue of $\mathbf{V}$ is $\omega_{k}=-\beta+\alpha\Theta_{k}$, where $\Theta_{k}$ denotes the $k$th eigenvalue of $\mathbf{D}$ [^16] If $\bm{z}_{k}$ denotes the $k$th eigenvector of $\mathbf{D}$, $\mathbf{D}\bm{z}_{k}=\Theta_{k}\bm{z}_{k}$ implies $\mathbf{V}\bm{z}_{k}=\omega_{k}\bm{z}_{k}$. Here, each $\omega_{k}$ measures the hypothetical utility gain for migrants if deviation occurs toward the $k$th eigenvector direction. In turn, $\omega^{*}=\max_{k}\{\omega_{k}\}$, and the stability of $\bar{\bm{x}}$ can be easily determined even in the many-region case. Observation 2 implies that such a clean approach is possible for all the models discussed earlier.
3.2   Maximum and minimum gains in proximity
It is thus important to understand the eigenvalue–eigenvector pairs of $\mathbf{D}$ before considering specific models. Once they are obtained, $\{\omega_{k}\}$ and $\omega^{*}$ are readily available. In fact, the eigenvalues $\{\Theta_{k}\}$ of $\mathbf{D}$ has a clear interpretation:
Observation 3.
@@ 3.2   Maximum and minimum gains in proximity
Each eigenvalue $\Theta_{k}$ of $\mathbf{D}$ measures the gain in proximity experienced by migrants when migration shocks occur in the direction of the corresponding eigenvector $\bm{z}_{k}$. ∎
To see this, consider a special case of the Beckmann model with $\beta=0$ and $\alpha=1$ where $\mathbf{V}=\mathbf{D}$ and $\omega_{k}=\Theta_{k}$. On the other hand, the indirect utility in this case is nothing but a parsimonious proximity measure $v_{i}(\bm{x})=\sum_{j\in\mathcal{I}}\phi_{ij}x_{j}$. This means that, each $\Theta_{k}$ ($=\omega_{k}$) represents the proximity gain for relocated agents when migration shock occurs toward the $\bm{z}_{k}$ direction.
As a concrete example, suppose $N=4$. The normalized proximity matrix is then
@@
There are two relevant eigenvalues: $\Theta_{1}=\frac{1-\phi}{1+\phi}$ and $\Theta_{2}=\big{(}\frac{1-\phi}{1+\phi}\big{)}^{2}$. An eigenvector associated with $\Theta_{1}$ is $\bm{z}_{1}=(1,0,-1,0)$, and represents a monocentric spatial pattern such as $\bar{\bm{x}}+\epsilon\bm{z}_{1}=(\bar{x},\bar{x}+\epsilon,\bar{x},\bar{x}-\epsilon)$. One region grows at the expense of another that is two steps away, creating a single center of attraction. In contrast, $\bm{z}_{2}=(1,-1,1,-1)$ associated with $\Theta_{2}$ corresponds to a polycentric agglomeration pattern $\bar{\bm{x}}+\epsilon\bm{z}_{2}=(\bar{x}+\epsilon,\bar{x}-\epsilon,\bar{x}+\epsilon,\bar{x}-\epsilon)$. Two regions located two steps apart grow symmetrically, while the two intermediate regions shrink.
There are three intuitive properties about $\Theta_{1}$ and $\Theta_{2}$. First, both are positive. Any deviation from the full dispersion induces some form of agglomeration, and thus increases the average proximity among movers. Second, each $\Theta_{k}$ decreases as $\phi$ increases. That is, the gain from proximity decreases if transport costs matter less. Third, $\Theta_{1}>\Theta_{2}$ at any value of $\phi$. Naturally, deviation toward a monocentric pattern induces a greater proximity gain than toward a polycentric pattern.
@@ D $k=8$
In fact, these properties generalize to the $N$-region case. [^17] Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2, provides the analytical formulae for $\{\Theta_{k}\}$ and $\{\bm{z}_{k}\}$, whereas the interpretation of $\{\Theta_{k}\}$ as proximity gains is newly given in the present study. Lemma 5 in Appendix A reproduces the relevant part of the aforementioned lemma, adapted to our context. In this case, each $\bm{z}_{k}$ corresponds to a deviation pattern with $k$ symmetric peaks, where $k=1,2,\ldots,\frac{N}{2}$, as illustrated in Fig. 6 for the $N=16$ case. Most importantly, the maximum and minimum gains in proximity are unambiguously determined:
Lemma 2.
Under Assumption C, $\max_{k}\Theta_{k}=\Theta_{1}$ and $\min_{k}\Theta_{k}=\Theta_{\frac{N}{2}}$ for all $\phi$. ∎
Proof.
See Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2. ∎
Intuitively, at any value of $\phi$, the monocentric agglomeration pattern ($k=1$) yields the largest proximity gain. The smallest gain arises from the pattern with the largest number of symmetric peaks, in which agents concentrate in every other region ($k=\frac{N}{2}$).
@@ 3.3   Contrasting implications for spatial patterns
Lemma 2 allows us to study the stability of the symmetric equilibrium, and to reveal the contrasting impact of local and global dispersion forces on spatial patterns.
In models with a local dispersion force, what matters is the maximum eigenvalue of $\mathbf{D}$, i.e., $\Theta_{1}$. For instance, in the Beckmann model, since $\omega_{k}=-\beta+\Theta_{k}$,
$\omega^{*}\gtreqless 0\quad\Leftrightarrow\quad\max_{k}\left\{-\beta+\alpha\Theta_{k}\right\}=-\beta+\alpha\max_{k}\left\{\Theta_{k}\right\}\gtreqless 0\quad\Leftrightarrow\quad\Theta_{1}\gtreqless\frac{\alpha}{\beta}.$ (24)
Figure 7A illustrates this for the $N=8$ case. The symmetry is stable if $\Theta_{1}$ is sufficiently small, i.e., if $\phi$ is sufficiently large. At the point where $\bar{\bm{x}}$ becomes unstable, a deviation of the form $\bar{\bm{x}}+\epsilon\bm{z}_{1}$ is self-reinforcing, and a single-peaked pattern as illustrated in Fig. 6A is favored by migrants.
@@ B The Braid model (Type G)
Figure 7: Curves of $\omega_{k}$ for the two minimal models ($N=8$, $k=1,2,3,4$).
Note: The uniform distribution is stable if all curves $\{\omega_{k}\}_{k=1}^{4}$ stay below the horizontal axis (gray region). It becomes unstable at $\phi^{*}$, at which the largest eigenvalue cuts the axis. In the Beckmann model, $\omega_{1}$ is the first to cross the axis, whereas in the Braid model, it is $\omega_{4}$.
In models with a global dispersion force, the minimum eigenvalue of $\mathbf{D}$ is critical. For example, in the Braid model, $\omega_{k}=\alpha\Theta_{k}-\Theta_{k}^{2}=\Theta_{k}(\alpha-\Theta_{k})$. Since $\Theta_{k}>0$, we see
$\omega^{*}\gtreqless 0\quad\Leftrightarrow\quad\max_{k}\left\{\alpha-\Theta_{k}\right\}=\alpha-\min_{k}\left\{\Theta_{k}\right\}\gtreqless 0\quad\Leftrightarrow\quad\alpha\gtreqless\Theta_{\frac{N}{2}},$ (25)
Figure 7B illustrates this for the $N=8$ case. The symmetry is stable if $\Theta_{\frac{N}{2}}$ is sufficiently large, that is, if $\phi$ is sufficiently small. At the point where $\bar{\bm{x}}$ becomes unstable, the $\frac{N}{2}$-centric deviation pattern $\bm{z}_{\frac{N}{2}}$ attracts agents, leading to the emergence of a polycentric pattern with the largest number of symmetric economic centers (Fig. 6D).
The minimal examples demonstrate that the preferred spatial pattern at the onset of instability of $\bar{\bm{x}}$ is formally different depending on the spatial scale of the dispersion force in a model. In particular, a polycentric spatial pattern is preferred only under a global dispersion force. This insight from the reduced-form models can be extended to cover general equilibrium models, such as those discussed in Section 2.5. We can summarize as follows:
@@
Consider a spatial model under Assumption C.
@@
    Consider a model of Type L or LG with local dispersion forces. Then, the uniform distribution $\bar{\bm{x}}$ can be stable at low transport cost levels. Suppose that the model parameters are set in a way that $\bar{\bm{x}}$ is stable for all $\phi\in(\phi^{*},1)$ with some $\phi^{*}\in(0,1)$, and becomes unstable at $\phi^{*}$, i.e., $\bar{\bm{x}}$ is unstable for $\phi$ slightly smaller than $\phi^{*}$. Then, a monocentric, single-peaked spatial equilibrium branches from $\bar{\bm{x}}$ at $\phi^{*}$.
@@
    Consider a model of Type G or LG with global dispersion forces. Then, $\bar{\bm{x}}$ can be stable at high transport cost levels. Suppose that the model parameters are set in a way that $\bar{\bm{x}}$ is stable for all $\phi\in(0,\phi^{*})$ with some $\phi^{*}\in(0,1)$, and becomes unstable at $\phi^{*}$, i.e., $\bar{\bm{x}}$ is unstable for $\phi$ slightly larger than $\phi^{*}$. Then, a polycentric spatial equilibrium with $\frac{N}{2}$ symmetric peaks branches from $\bar{\bm{x}}$ at $\phi^{*}$.
Proof.
See Section A.2. ∎
Note that Proposition 1 focuses on the case where (i) multiple equilibria are possible but at the same time (ii) $\bar{\bm{x}}$ is stable for some $\phi$. If local dispersion forces are sufficiently strong in a Type L (or LG) model, $\bar{\bm{x}}$ is stable for all $\phi\in(0,1)$, and $\phi^{*}$ mentioned in Proposition 1 (b) may not exist. In this case, the equilibrium is unique for all transport cost levels. This convenient property has allowed unambiguous counterfactual analysis in conventional quantitative spatial models. It is noted that, however, uniqueness of equilibrium implies that $\bar{\bm{x}}$ is the only possible outcome under the absence of exogenous geographical asymmetries. Also, for Type G models, $\bar{\bm{x}}$ can be unstable for all $\phi$ if agglomeration force is too strong. In this case, all agents may concentrate in one location. Proposition 1 excludes these cases so that the level of transport costs matters.

### v8

@@ 3.1  The stability of the symmetric equilibrium
For any deviation pattern $\bm{z}$ with $\sum_{i\in\mathcal{I}}z_{i}=0$ and $\|\bm{z}\|=1$, the expected utility gain of marginal movers is $\omega(\bm{z})\equiv\bm{z}^{\top}\mathbf{V}\bm{z}$, where $\top$ denotes transpose. ∎
Then, the value $\omega(\bm{z})$ characterizes the stability of $\bar{\bm{x}}$. If $\omega(\bm{z})<0$ for all possible deviation patterns, then any migration shock reduces the utility of the movers and $\bar{\bm{x}}$ is locally stable. If $\omega(\bm{z})>0$ for some $\bm{z}$, the movers benefit from relocating according to that pattern, implying instability. Thus, stability is determined by whether the maximal attainable gain $\omega^{*}\equiv\max_{\bm{z}}\omega(\bm{z})$ is positive or negative.
To characterize $\omega^{*}$, let $\{\omega_{k}\}$ denote the eigenvalues of $\mathbf{V}$, and $\{\bm{z}_{k}\}$ be their associated eigenvectors. Because $\omega(\bm{z}_{k})=\bm{z}_{k}^{\top}\mathbf{V}\bm{z}_{k}=\omega_{k}\bm{z}_{k}^{\top}\bm{z}_{k}=\omega_{k}\|\bm{z}_{k}\|^{2}=\omega_{k}$, we observe
$\omega^{*}=\max_{k}\{\omega_{k}\},$ (32)
implying that the largest eigenvalue of $\mathbf{V}$ determines the stability of $\bar{\bm{x}}$. [^15] The discussion here is closely related to Allen-etal-AER2024. In their framework, the spectral radius of a matrix that collects key model elasticities governs the uniqueness of the equilibrium. In our setting, their uniqueness condition [Theorem 1(a)] broadly corresponds to a sufficient condition for the stability of $\bar{\bm{x}}$ for all $\phi$, which rules out endogenous agglomeration. Our focus instead lies on environments with multiple equilibria [cf. their Theorem 1(c)], and we take a step toward understanding how the underlying network structure shapes the positive properties of these equilibria.
How can we obtain $\{\omega_{k}\}$? Under Assumption˜C, for all models in Table˜1, the benefit matrix satisfies $\mathbf{V}=\Omega(\mathbf{D})$, where the row-normalized proximity matrix $\mathbf{D}$ is replaced by its $N\times N$ counterpart. For each model, the base gain function $\Omega(\Theta)$ is the same as in the case of two-regions. [^16] This is simply because, for all models we saw in Section 2.4, the gain functions are derived for the general number of locations under Assumption C, and then specialized to the $N=2$ case. Then, $\mathbf{V}$ and $\mathbf{D}$ share the same set of eigenvectors $\{\bm{z}_{k}\}$, and $\mathbf{V}=\Omega(\mathbf{D})$ implies
$\omega_{k}=\Omega(\Theta_{k}),$ (33)
where $\Theta_{k}$ is the eigenvalue of $\mathbf{D}$ corresponding to $\bm{z}_{k}$. Each $\Theta_{k}$ is a function of $\phi$ because $\phi$ is the only parameter of $\mathbf{D}$ according to our assumptions.
These observations yield a transparent approach to stability analysis, as the following example illustrates.
Example 3.
In the Beckmann model, we again have $\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D}$, where $\mathbf{I}$ and $\mathbf{D}$ are replaced by the $N$-region identity and row-normalized proximity matrices. [^17] Observe that Eq. 2 is defined for general $N$. The $k$th eigenvalue of $\mathbf{V}$ is then $\omega_{k}=-\beta+\alpha\Theta_{k}$ because $\mathbf{D}\bm{z}_{k}=\Theta_{k}\bm{z}_{k}$ implies $\mathbf{V}\bm{z}_{k}=(-\beta+\alpha\Theta_{k})\bm{z}_{k}$. The stability of $\bar{\bm{x}}$ is determined by the sign of $\omega^{*}=\max_{k}\{\omega_{k}\}$. Consider a value of $\phi$ such that $\bar{\bm{x}}$ is stable, i.e., $\omega_{k}<0$ for all $k$ and hence $\omega^{*}=\max_{k}\{\omega_{k}\}<0$. Suppose then that $\phi$ increases or decreases monotonically. Suppose some $\omega_{k*}$ changes its sign from negative to positive at some $\phi^{*}$. Then, after that point, deviation towards the $\bm{z}_{k^{*}}$-direction improves the utility of the movers, and $\bar{\bm{x}}$ becomes unstable. In this sense, the corresponding eigenvector $\bm{z}_{k^{*}}$ represents the critical deviation pattern. ∎
3.2  Proximity gains and deviation patterns
The eigenvalues $\{\Theta_{k}\}$ of $\mathbf{D}$ have a clear interpretation analogous to the proximity gain $\Theta$ in the two-region case. Again, for the special case of the Beckmann model with $\beta=0$ and $\alpha=1$, the indirect utility function boils down to a simple proximity measure. We have $\mathbf{V}=\mathbf{D}$ and $\omega_{k}=\Theta_{k}$, which can be interpreted as follows.
Observation 1.
@@ 3.2  Proximity gains and deviation patterns
Each eigenvalue $\Theta_{k}$ of $\mathbf{D}$ measures the proximity gain experienced by marginal movers when migration shocks occur in the corresponding direction, $\bm{z}_{k}$. ∎
For example, if $N=4$, the normalized proximity matrix under Assumption˜C is
@@
There are two relevant eigenvalues:
$\Theta_{1}=\frac{1-\phi}{1+\phi}\quad\text{and}\quad\Theta_{2}=\left(\frac{1-\phi}{1+\phi}\right)^{2}.$ (35)
We can check that there are two eigenvectors associated with $\Theta_{1}$, namely $\bm{z}_{1}^{+}=(1,0,-1,0)$ and $\bm{z}_{1}^{-}=(0,1,0,-1)$. Both represent monocentric spatial patterns. For example, in $\bar{\bm{x}}+\epsilon\bm{z}_{1}^{-}=(\bar{x},\bar{x}+\epsilon,\bar{x},\bar{x}-\epsilon)$, one region grows at the expense of another that is two steps away, creating a single center of attraction. For $\Theta_{2}$, the associated eigenvector is $\bm{z}_{2}=(1,-1,1,-1)$, and represents a polycentric agglomeration pattern $\bar{\bm{x}}+\epsilon\bm{z}_{2}=(\bar{x}+\epsilon,\bar{x}-\epsilon,\bar{x}+\epsilon,\bar{x}-\epsilon)$. Two regions located two steps apart grow symmetrically, while the two intermediate regions shrink.
There are three intuitive properties about $\Theta_{1}$ and $\Theta_{2}$. First, both are positive. Any deviation from the full dispersion induces some form of agglomeration, and for movers this increases the proximity to others. Second, each $\Theta_{k}$ decreases as $\phi$ increases. The proximity gain decreases if transport costs are less important. Third, $\Theta_{1}>\Theta_{2}$ at any value of $\phi$. Naturally, from the perspective of movers, deviation toward a monocentric pattern induces a greater proximity gain than toward a polycentric pattern.
@@ D $k=8$
These properties generalize to the $N$-region case. [^18] Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2, provides the analytical formulae for $\{\Theta_{k}\}$ and $\{\bm{z}_{k}\}$, while the interpretation of $\{\Theta_{k}\}$ as proximity gains is newly given in the present study. Lemma 5 in Appendix A reproduces the relevant part of the aforementioned lemma, adapted to our context. All $\{\Theta_{k}\}$ are positive and decrease in $\phi$. Each $\bm{z}_{k}$ corresponds to a deviation pattern with $k$ symmetric peaks as illustrated in Fig.˜7. Most importantly, the maximum and minimum proximity gains are unambiguously determined: for each $\phi$,
$\max_{k}\Theta_{k}=\Theta_{1}\quad\text{and}\quad\min_{k}\Theta_{k}=\Theta_{\frac{N}{2}}.$ (36)
Intuitively, at any value of $\phi$, the monocentric agglomeration pattern ($k=1$, Fig.˜7A) yields the largest proximity gain for the movers. The proximity gain is the smallest if the movers agglomerates in every other region ($k=\frac{N}{2}$, Fig.˜7D).
@@ 3.3  Contrasting implications for spatial patterns
The relationship (36) has important implications for endogenous spatial patterns. This is because the maximum eigenvalue of $\mathbf{D}$ is critical if the model incorporates only local dispersion forces. For instance, since $\omega_{k}=-\beta+\Theta_{k}$ in the Beckmann model, we see
$\omega^{*}<0\quad\Leftrightarrow\quad\max_{k}\left\{-\beta+\alpha\Theta_{k}\right\}=-\beta+\alpha\max_{k}\left\{\Theta_{k}\right\}<0\quad\Leftrightarrow\quad\Theta_{1}<\frac{\alpha}{\beta}.$ (37)
Figure˜8A illustrates this for the $N=8$ case, where we draw curves of $\omega_{k}$ as a function of $\phi$. The symmetric equilibrium is stable if $\Theta_{1}$ is sufficiently small, i.e., if $\phi$ is sufficiently large. If $\phi$ monotonically decreases from a high level and crosses $\phi^{*}$, then $\bar{\bm{x}}$ becomes unstable at $\phi^{*}$ because a deviation of the form $\bar{\bm{x}}+\epsilon\bm{z}_{1}$ induces a positive utility gain for the movers. Monocentric agglomeration should form at such a point.
@@ B The Braid model (Type G)
Figure 8: Curves of $\omega_{k}=\Omega(\Theta_{k})$ for the two minimal models ($N=8$, $k=1,2,3,4$).
Note: The symmetric equilibrium is stable if all curves $\{\omega_{k}\}_{k=1}^{4}$ stay below the horizontal axis (gray region). It becomes unstable at $\phi^{*}$ where the largest eigenvalue cuts the axis. In the Beckmann model, $\omega_{1}$ is the first to cross the axis, whereas in the Braid model, it is $\omega_{4}$.
By contrast, if the model has only global dispersion forces, the minimum eigenvalue of $\mathbf{D}$ is critical. In the Braid model, $\omega_{k}=\alpha\Theta_{k}-\Theta_{k}^{2}=\Theta_{k}(\alpha-\Theta_{k})$, which implies that
$\omega^{*}<0\quad\Leftrightarrow\quad\max_{k}\left\{\alpha-\Theta_{k}\right\}=\alpha-\min_{k}\left\{\Theta_{k}\right\}<0\quad\Leftrightarrow\quad\alpha<\Theta_{\frac{N}{2}}.$ (38)
Figure˜8B illustrates this for the $N=8$ case. The symmetric equilibrium is stable if $\Theta_{\frac{N}{2}}$ is sufficiently large, that is, if $\phi$ is sufficiently small. If $\phi$ gradually increases from a very small value, at $\phi^{*}$, the $\frac{N}{2}$-centric deviation pattern $\bm{z}_{\frac{N}{2}}$ becomes attractive for movers.
The two minimal examples show that the spatial pattern emerging at the onset of instability is critically dependent on the spatial scale of the dispersion force in the model. In particular, polycentric patterns arise only when the dispersion force operates at a global scale. This insight from the reduced-form models can be extended to cover general equilibrium models discussed in Section˜2.4:
@@
Suppose Assumption˜C.
@@
    Consider a model of Type L or LG with local dispersion forces. Then, the symmetric equilibrium $\bar{\bm{x}}$ is stable for large $\phi$. Suppose that the model parameters are set so that $\bar{\bm{x}}$ is stable for all $\phi>\phi^{*}$ with some threshold value $\phi^{*}\in(0,1)$, and becomes unstable at $\phi^{*}$, i.e., $\bar{\bm{x}}$ is unstable for $\phi$ slightly smaller than $\phi^{*}$. Then, a single-peaked monocentric spatial equilibrium path branches from $\bar{\bm{x}}$ at $\phi^{*}$.
@@
    Consider a model of Type G or LG with global dispersion forces. Then, $\bar{\bm{x}}$ can be stable for small $\phi$. Suppose that the model parameters are set so that $\bar{\bm{x}}$ is stable for all $\phi<\phi^{*}$ with some threshold value $\phi^{*}\in(0,1)$, and becomes unstable at $\phi^{*}$, i.e., $\bar{\bm{x}}$ is unstable for $\phi$ slightly larger than $\phi^{*}$. Then, a polycentric spatial equilibrium path with $\frac{N}{2}$ symmetric peaks branches from $\bar{\bm{x}}$ at $\phi^{*}$.
Proposition˜1 considers settings in which multiple equilibria may arise, while $\bar{\bm{x}}$ remains stable for some values of $\phi$. This condition need not always hold. For example, in a Type L model with sufficiently strong local dispersion forces, $\bar{\bm{x}}$ is stable for all $\phi\in(0,1)$, and the threshold $\phi^{*}$ in Proposition˜1 (b) does not exist. In such cases, the equilibrium is typically unique, a convenient feature that makes unambiguous counterfactual analysis feasible in quantitative spatial models. Equilibrium uniqueness implies that, absent exogenous geographical asymmetries, $\bar{\bm{x}}$ is the only possible outcome. At the opposite extreme, in Type G models with sufficiently strong agglomeration forces, $\bar{\bm{x}}$ can be unstable for all $\phi$, leading all agents to concentrate in a single location. Proposition˜1 deliberately excludes both of these extremal cases.